Stinehelfer Beet Processors, Inc., processes sugar beets in batches. A batch of sugar beets costs $52 to buy from farmers and $14 to crush in the company's plant. Two intermediate products, beet fiber and beet juice, emerge from the crushing process. The beet fiber can be sold as is for $28 or processed further for $20 to make the end product industrial fiber that is sold for $40. The beet juice can be sold as is for $48 or processed further for $32 to make the end product refined sugar that is sold for $74. What is the financial advantage (disadvantage) for the company from processing the intermediate product beet juice into refined sugar rather than selling it as is?

Garrison 16e Rechecks 2017-12-15

Multiple Choice

  • ($6)

  • $8

  • ($35)

  • $60

Swagger Corporation purchases potatoes from farmers. The potatoes are then peeled, producing two intermediate products-peels and depeeled spuds. The peels can then be processed further to make a cocktail of organic nutrients. And the depeeled spuds can be processed further to make frozen french fries. A batch of potatoes costs $46 to buy from farmers and $12 to peel in the company's plant. The peels produced from a batch can be sold as is for animal feed for $28 or processed further for $17 to make the cocktail of nutrients that are sold for $48. The depeeled spuds can be sold as is for $39 or processed further for $28 to make frozen french fries that are sold for $60.

Required:

Assuming that no other costs are involved in processing potatoes or in selling products, how much money does the company make from processing one batch of potatoes into the cocktail of organic nutrients and frozen french fries?

Should each of the intermediate products, peels and depeeled spuds, be sold as is or processed further into an end product?

Cocktail of Organic Nutrients Frozen French Fries
Financial advantage (disadvantage) from further processing
Sold as is or processed further?
Answer #1

Financial Advantage = Selling price after processing – Selling price before processing – Processing costs

Joint costs incurred upto the split off point are sunk costs and not relevant

Hence, advantage = 74 – 48-32

= ($6)

Money from one batch = 48+60-28-17-46-12

= $5 per batch


Cocktail of Organic Nutrients

Frozen French Fries

Financial advantage (disadvantage) from further processing

48-28-17 = 3

60-39-28 = ($7)

Sold as is or processed further?

Process further

Sell as is